OpenStudy (anonymous):

Suppose a particular jet needs to attain a speed of 220 mph to take off. If it can accelerate from 0 to 220 mph in 45 seconds, how long must the runway be? Assume constant acceleration. Note: 1 mph = 22/15 ft/sec. (answer in feet)

OpenStudy (anonymous):

a=(v-u)/t put v=225 and u=0 and t=45 we can then get a by s=ut+(1/2)(a)t^2 put u=0, t=45 and a=what we get from the previous step we can then get the value of s which is distance

OpenStudy (anonymous):

distance in miles right?

OpenStudy (anonymous):

7260 ft right?